# Short proof of the hypergraph container theorem

### v1

@@ 1 Introduction
Let $V$ be a finite set. Given a subset $X\subseteq V$, let $\langle X\rangle=\{S\subseteq V\colon X\subseteq S\}$. We say a probability measure $\nu$ over $2^{V}$ is $(p,K)$-uniformly-spread if for every $X\subseteq V$ we have $\nu(\langle X\rangle)\leq Kp^{|X|-1}/|V|$. Uniform signifies that the measure is fairly uniform from the point of view of elements of $V$. Throughout the paper we use $V=V(\mathcal{H})$ and $N=|V|$, where $\mathcal{H}$ is a given hypegraph.
@@
For every $\ell\in\mathbb{N}$ and $K,\varepsilon>0$ there exists $T>0$ such that the following holds. Suppose $\mathcal{H}$ is an $\ell$-uniform hypergraph ($\ell$-graph for short), and let $\nu$ be $(p,K)$-uniformly-spread measure over $2^{V}$, supported on $\mathcal{H}$. Then for every independent set $I\subseteq V(\mathcal{H})$ there exists $F\subseteq I$ and $C=C(F)\subseteq V$ such that $|F|\leq TNp$, $\nu(\mathcal{H}[C])<\varepsilon$, and $I\subseteq C\cup F$.
@@ 2 Proof
For every $\ell\in\mathbb{N}$ and $K>0$ there exists $\delta>0$ such that the following holds. Suppose $\mathcal{H}$ is an $\ell$-graph, and let $\nu$ be a $(p,K)$-uniformly-spread measure over $2^{V}$, supported on $\mathcal{H}$. Then for every independent set $I\subseteq V$ there exists $F\subseteq I$ and $C=C(F)\subseteq V$ such that $|F|\leq\ell Np$, $|C|\leq(1-\delta)N$, and $I\subseteq C\cup F$. Moreover, $C$ can be unambigously constructed from any $F\subseteq\hat{F}\subseteq I$.
@@
Let $\alpha=2^{-\ell-2}$. We now distinguish two cases, where if $\nu(\mathcal{H}^{\prime})$ is large, then we can apply the inductive hypothesis to the $(\ell-1)$-uniform hypergraph consisting of the $(\ell-1)$-shadow of $\mathcal{H}^{\prime}$, and otherwise we can immediately find a small container $C$ for which $I\smallsetminus F\subseteq C$.
Case 1: $\nu(\mathcal{H}^{\prime})\geq\alpha p$. Set $\nu^{\prime}$ to be the probability measure over $2^{V}$ given by
$\nu^{\prime}(X)\,\propto\,\begin{cases}\nu(\langle X\rangle\cap\mathcal{H}^{\prime}),&\text{if }|X|=\ell-1,\\ 0,&\text{otherwise.}\end{cases}$
From (2) and $\nu(\mathcal{H}^{\prime})\geq\alpha p$ we conclude that $\nu^{\prime}$ is $(2K\alpha^{-1},p)$-uniformly-spread. Let $\mathcal{H}^{\prime\prime}$ be the $(\ell-1)$-graph given by the support of $\nu^{\prime}$. Then $I$ is also an independent set in $\mathcal{H}^{\prime\prime}$, thus by the induction hypothesis there exists $F^{\prime}\subseteq V$ of size $|F^{\prime}|\leq(\ell-1)Np$ and $C=C(F^{\prime})$ such that $|C|\leq(1-\delta)N$ and $I\subseteq C\cup F^{\prime}$. Note that we can reconstruct $C$ from $F:=F\cup F^{\prime}$.
@@
For the sake of completeness, we derive Theorem 1.1 from Lemma 2.1 implies.
@@
Let $\delta>0$ be as given by Lemma 2.1 for $\ell$ and $K^{2}\varepsilon^{-2}$ (as $K$). We prove the lemma for $T=\ell\log(K\varepsilon^{-1})/\log(1+\delta)$.
@@
If $\nu(\mathcal{H}[C])\geq\varepsilon$ then $|C|\geq\varepsilon N/K$. Therefore, $\nu^{\prime}$ is $(p,K^{2}\varepsilon^{-2})$-uniformly-spread each time we apply Lemma 2.1. In each iteration the set $C$ shrinks by a factor of $1-\delta$, thus we are done after at most $\log(K\varepsilon^{-1})/\log(1+\delta)$ iterations. The set $F$ grows by at most $\ell Np$ in each iteration, which gives an upper bound of $TNp$ on its final size for appropriately chosen $T=T(K,\varepsilon)$. Due to the last property in Lemma 2.1, the final set $C$ can be unambiguously constructed from $F$. ∎
@@ Acknowledgment.
Ideas used in this paper were developed while the first author was visiting Stanford University in November 2023. The first author thanks Jacob Fox for hospitality.

### v3

@@ 1 Introduction
Let $V$ be a finite set. Given a subset $X\subseteq V$, let $\langle X\rangle=\{S\subseteq V\colon X\subseteq S\}$. We say a probability measure $\nu$ over $2^{V}$ is $(p,K)$-uniformly-spread if for every non-empty $X\subseteq V$ we have $\nu(\langle X\rangle)\leq Kp^{|X|-1}/|V|$. Uniform signifies that the measure is fairly uniform from the point of view of elements of $V$. Throughout the paper we use $V=V(\mathcal{H})$ and $N=|V|$, where $\mathcal{H}$ is a given hypergraph. If all edges in a hypergraph $\mathcal{H}$ have size at most $\ell$, we say that $\mathcal{H}$ is an $(\leq\ell)$-graph.
@@
For every $\ell\in\mathbb{N}$ and $K,\varepsilon>0$ there exists $T>0$ such that the following holds. Suppose $\mathcal{H}$ is an $(\leq\ell)$-graph, and let $\nu$ be $(p,K)$-uniformly-spread measure over $2^{V}$ supported on $\mathcal{H}$, for some $p\in(0,1]$. Then for every independent set $I\subseteq V(\mathcal{H})$ there exists $F\subseteq I$ and $C=C(F)\subseteq V$ such that $|F|\leq TNp$, $\nu(\mathcal{H}[C])<\varepsilon$, and $I\subseteq C\cup F$.
@@ 2 Proof
For every $\ell\in\mathbb{N}$ and $K>0$ there exists $\delta>0$ such that the following holds. Suppose $\mathcal{H}$ is an $(\leq\ell)$-graph, and let $\nu$ be a $(p,K)$-uniformly-spread measure over $2^{V}$ supported on $\mathcal{H}$, for some $p\in(0,1]$. Then for every independent set $I\subseteq V$ there exists $F\subseteq I$ and $C=C(F)\subseteq V$ such that $|F|\leq\ell Np$, $|C|\leq(1-\delta)N$, and $I\subseteq C\cup F$. Moreover, $C$ can be unambigously constructed from any $F\subseteq\hat{F}\subseteq I$.
@@
Let $\alpha=2^{-\ell-2}$. We now distinguish two cases, where if $\nu(\mathcal{H}^{\prime})$ is large, then we can apply the inductive hypothesis to an appropriate $(\leq\ell-1)$-graph, and otherwise we can immediately find a small container $C$ for which $I\smallsetminus F\subseteq C$.
Case 1: $\nu(\mathcal{H}^{\prime})\geq\alpha p$. Let $\mathcal{H}^{\prime\prime}$ denote the $(\leq\ell-1)$-graph consisting of sets $X$ such that $X=H^{\prime}\smallsetminus F$ for some $H^{\prime}\in\mathcal{H}^{\prime}$. Set $\nu^{\prime}$ to be the probability measure over $2^{V\smallsetminus F}$ given by
$\nu^{\prime}(X)\,\propto\,\begin{cases}\nu((X\cup 2^{F})\cap\mathcal{H}^{\prime}),&\text{if }X\in\mathcal{H}^{\prime\prime},\\ 0,&\text{otherwise.}\end{cases}$
From (2) and $\nu(\mathcal{H}^{\prime})\geq\alpha p$ we conclude that $\nu^{\prime}$ is $(2K\alpha^{-1},p)$-uniformly-spread. Also observe that $I$ is an independent set in $\mathcal{H}^{\prime\prime}$, thus by the induction hypothesis there exists $F^{\prime}\subseteq V$ of size $|F^{\prime}|\leq(\ell-1)Np$ and $C=C(F^{\prime})$ such that $|C|\leq(1-\delta)N$ and $I\subseteq C\cup F^{\prime}$. Note that we can reconstruct $C$ from $F:=F\cup F^{\prime}$.
@@
For the sake of completeness, we derive Theorem 1.1 from Lemma 2.1.
@@
Let $\delta>0$ be as given by Lemma 2.1 for $\ell$ and $K/\varepsilon$ (as $K$). We prove the theorem for $T=\ell\log(K\varepsilon^{-1})/\log(1+\delta)$.
@@
If $\nu(\mathcal{H}[C])\geq\varepsilon$, then for nonempty $X\subseteq C$,
$\nu^{\prime}(\langle X\rangle)\leq\frac{\nu(\langle X\rangle)}{\nu(\mathcal{H}[C])}\leq\frac{Kp^{|X|-1}/N}{\varepsilon}\leq\frac{K}{\varepsilon}p^{|X|-1}/|C|,$
and hence $\nu^{\prime}$ is $(p,K/\varepsilon)$-uniformly-spread each time we apply Lemma 2.1. Furthermore, if $\nu(\mathcal{H}[C])\geq\varepsilon$, then $|C|\geq\varepsilon N/K$. In each iteration the set $C$ shrinks by a factor of $1-\delta$, thus we are done after at most $\log(K\varepsilon^{-1})/\log(1+\delta)$ iterations. The set $F$ grows by at most $\ell Np$ in each iteration, which gives an upper bound of $TNp$ on its final size for the above choice of $T=T(K,\varepsilon)$. Due to the last property in Lemma 2.1, the final set $C$ can be unambiguously constructed from $F$. ∎
@@ Acknowledgment.
Ideas used in this paper were developed while the first author was visiting Stanford University in November 2023. The first author thanks Jacob Fox for hospitality. We thank Jacob Fox for helpful comments and Wojciech Samotij for pointing out a subtle issue in an earlier version.